Grade 12Algebra

(1+x)n=C0+C1x+....Cn
where Cr=nCr & (C0 + C1)(C1 + C2)
(Cn-1 + Cn) = AC1C2.....Cn
when n=5
Find A.

Purva
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
Dear student
 
 
we know the formula:
nCr + nC(r-1) = (n+1)Cr
so, (C0 + C1)(C+ C2).....(Cn-1 + Cn) = 6C1*6C2*...*6C5
now you can find A simply by calculating 6C1*6C2*...*6C5/5C1*5C2*...*5C5
 
 
Hope it helps
 
Regards
Arun (askIITians forum expert)